RC meeting and Run Plan, Sunday December 12th
RC Update
Run Coordinator: Brian Quinn
No RC meeting this weekend
Upcoming Shifts Which Need Filled
A total of 3 vacant shifts remain until the winter break shutdown:
- Wed Dec 15th, OWL, SL
- Fri Dec 17th, OWL, TO
- Sat Dec 18th, OWL, TO
Thanks!!! to Thir, Scott, Nathaniel, Maria, Bob, Sean, Albert, and Zeke!!!... and especially Kondo who filled in for Saturday day, at short notice, allowing us to keep running!!
Accelerator and Hall Status
- 4-hour optics studies 7 AM Tuesday now seems to be certain.
- Requested controlled access.
- There may be some delay getting sweep, all halls will probably go to controlled.
Special runs done
- Friday Day: low current (2 uA) test runs on LD2 with GEM HVs decreased by 50 V, 25 V, and 0V, relative to nominal.
- Friday Swing: SBS-off running 2hr LH2, 3hr LD2 (With SBS-off HV-settings for BBCal)
- Saturday Day: BBCal cosmics 3.2 M events (run 12744)
- Saturday Swing: GRINCH-LED-only runs at 100 Hz with varied HV
- Sunday Day: HCal-LED-only (run 12775)
Unfortunately, the last 3 were made possible by long accelerator-down periods... including one caused by bird-droppings in Richmond!
To be Done During Extended Down Time
- Removal of INFN J2 from Hall. (Jessie) Zeke would like it to be ASAP after end of December run.
Run Plan
- Finished Cycle 14 today
- Now taking LD2 data for Cycle 15
- Continue following cycle pattern and ticking off runs-completed on paper copy at SL desk
(Reviewing the Run List shows that the check-off list has gotten out of phase with the numbering of the cycles. Rather than re-numbering the cycles we've been taking, I annotated the Run List with 'long cycles' (20 M LH2, 80 M LD2) starting with Cycle -3! That's followed by cycles -2, -1, 0 and 1. Using that numbering system, we are now on Cycle 15. There were also 7 'short' cycles (10 M LH2, 40 M LD2). )
Opportunistic Access Requests for Tuesday
- Put network-readable 'scope on trigger signals to quickly identify 70 ns timing shifts. (Alexandre)
- Try changing HV channel for J0 lower module (to see whether lower current is a result of bad read-back) and perhaps check HV divider (if accessible) (Zeke or Holly).
- Remove unused HV card from HV mainframe for hodoscope (Ralph)
- GRINCH (change pulser brightness) (Todd/Maria)
- Maybe, if rate of failure increases and replacement rpi and be prepared in time, change rpi & SD card in HCal's upper HV mainframe (Sebastian & Ellen)
Resolved Issues
- Low DAQ rate alarm for hung MPD (End, Reset(twice?), configure, download)
- GRINCH LED pulser now prescaled (PS5=4) to give ~10 Hz
- "MPD NotFound" error:
- See two-page printout from GEM documentation, on SL desk
- On page 3, section 3.3 (marked by bracket in the margin) tells how to boot the VTP crate
- If that doesn't fix it, see the next page for how to boot UVa MPD crate (or call Alexandre or GEM expert on call)
Unresolved Issues
- Hodoscope HV alarms (Ralph and Steve are working on it)
- BBCal HV alarms with CH HV-control showing HV off by more than 50 V while JavaGUI shows it at setpoint
- BBCal HV alarm with JavaGui showing HV at ~500 V
- HCal HV changes (contact Scott or Sebastian to reload the values and make sure everything get's properly loaded)
- Occasional timing shifts of ~70 ns
- GRINCH sync errors (can be ignored)